# Dept. of Systems Engineering, ETS Montreal, Canada

### v1

@@ 5 Conclusion
This paper introduces the Bidirectional Multi-step Domain Generalization (BMDG) framework for V-I ReID. It uses a novel prototype learning approach that learns the most discriminative and complementary set of prototypes from both modalities. This set of prototypes generates multiple intermediate domains between modalities by progressively mixing prototypes from each modality, reducing the domain gap. The effectiveness of BMDG is shown experimentally on the SYSU-MM01 and RegDB datasets, where it outperforms state-of-the-art methods for V-I person ReID. Results also highlight the impact of considering multiple intermediate steps and bidirectional training to improve the model’s accuracy significantly. Moreover, BMDG can be applied to other state-of-the-art part-based V-I ReID methods, significantly improving their ReID accuracy.
Appendix 0.A Proofs
@@
0.A.1 Maximizing $\text{MI}(P^{1},\dots,P^{K};Y)$
@@
$\text{MI}(P^{1};Y)+\dots+\text{MI}(P^{K};Y),$ (0.A.1)
@@
$\text{MI}(X,Z;Y)=\text{MI}(X;Y)+\text{MI}(Z;Y)-\text{MI}(X;Z;Y)$ (0.A.5)
$MI(P^{1},\dots,P^{K};Y).\vspace{-.2cm}$ (0.A.6)
Figure 0.A.1: Venn diagram of theoretic measures for three variables $X$, $Y$, and $Z$, represented by the lower left, upper, and lower right circles, respectively.
Theorem 0.A.1
[…]
Proof
@@
$\text{MI}(P^{1},\tilde{P}^{1};Y)=\underbrace{\text{MI}(P^{1};Y)}_{\alpha}+\underbrace{\text{MI}(\tilde{P}^{1};Y)}_{\beta}-\underbrace{\text{MI}(P^{1};\tilde{P}^{1};Y)}_{\gamma},$ (0.A.7)
@@
$\text{MI}(P^{1};\tilde{P}^{1};Y)\leq\text{MI}(P^{1};\tilde{P}^{1}),$ (0.A.8)
So by minimizing the right term of Eq. 0.A.8, $\gamma$ is also minimized. We show that $\text{MI}(P^{1};\tilde{P}^{1})=0$ by expanding $\tilde{P}^{1}$ to $(P^{2},\tilde{P}^{2})$:
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=\text{MI}(P^{1};P^{2})+\text{MI}(P^{1};\tilde{P}^{2})-\text{MI}(P^{1};P^{2};\tilde{P}^{2}).$ (0.A.9)
@@
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=\text{MI}(P^{1};\tilde{P}^{2}).$ (0.A.10)
@@
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=0.$ (0.A.11)
To maximize $\beta$, we can rewrite and expand recursively in Eq. 0.A.7 :
$\text{MI}(\tilde{P}^{1};Y)=\text{MI}(P^{2},\tilde{P}^{2};Y)=\underbrace{\text{MI}(P^{2};Y)}_{\text{maximizing}}+\underbrace{\text{MI}(\tilde{P}^{2};Y)}_{\text{expanding}}-\underbrace{\text{MI}(P^{2};\tilde{P}^{2};Y)}_{0}.$ (0.A.12)
@@
$\text{MI}(\tilde{P}^{k};Y)=\text{MI}(P^{k+1},\tilde{P}^{k+1};Y)=\text{MI}(P^{k};Y)+\dots+\text{MI}(P^{K};Y)\;\;\;\forall k\in\{0,\dots,K-1\}.$ (0.A.13)
@@
$\text{MI}(P^{1},\dots,P^{K};Y)=\sum_{k=1}^{K}{\text{MI}(P^{k};Y)},$ (0.A.14)
@@
0.A.2 Maximizing $\text{MI}(P^{k};Y)$
In Section 3, the MI between the representation of each prototype and the label of persons are maximized by minimizing cross-entropy loss (see Eq: 11 of the manuscript). This approximation is formulated as Proposition 1.
@@ 0.A.2 Maximizing $\text{MI}(P^{k};Y)$
[…]
Proof
@@
$\text{MI}(P^{k};Y)=\underbrace{\mathcal{H}(Y)}_{\delta}-\underbrace{\mathcal{H}(Y|P^{k})}_{\xi}$ (0.A.15)
@@
$\mathcal{H}(Y;\hat{Y}|P^{k})=\mathcal{H}(Y|P^{k})+\underbrace{\mathcal{D}_{\text{KL}}(Y||\hat{Y}|P^{k})}_{\geq 0},$ (0.A.16)
@@
$\mathcal{H}(Y|P^{k})\leq\mathcal{H}(Y;\hat{Y}|P^{k}).$ (0.A.17)
Through the minimization of Eq. 11, training can naturally be decoupled in 2 steps. First, weights of the prototype module are fixed, and only the classifier parameters (i.e., weight $W^{k}$ of the fully connected layer) are minimized w.r.t. Eq. 0.A.16. Through this step, $\mathcal{D}_{\text{KL}}(Y||\hat{Y}|P^{k})$ is minimized by adjusting $\hat{Y}$ while the $\mathcal{H}(Y|P^{k})$ does not change. In the second step, the prototype module’s weights are minimized w.r.t. $\mathcal{H}(Y|P^{k})$, while the classifier parameters $W^{k}$ are fixed.
Appendix 0.B Training Algorithm
@@
Appendix 0.C Additional Details on the Experimental Methodology
0.C.1 Datasets:
Research on cross-modal V-I ReID has extensively used the SYSU-MM01 [SYSU] and RegDB [regDB] datasets. SYSU-MM01 is a large dataset containing more than 22K RGB and 11K IR images of 491 individuals captured with 4 RGB and 2 near-IR cameras, respectively. Of the 491 identities, 395 were dedicated to training, and 96 were dedicated to testing. Depending on the number of images in the gallery, the dataset has two evaluation modes: single-shot and multi-shot. RegDB contains 4,120 co-located V-I images of 412 individuals. Ten trial configurations randomly divide the dataset into two sets of 206 identities for training and testing. The tests are conducted in two ways – comparing I to V (query) and vice versa.
0.C.2 Experimental protocol:
[…]
0.C.3 Performance measures:
@@
Appendix 0.D Additional Quantitative Results
0.D.1 Hyperparameter values:
[…]
0.D.2 Step size and number of part prototypes:
In section 4.2, we discussed the step size and number of part prototypes based on Rank-1 accuracy. Here we report the mAP measurement for Table 2(a) in paper in Table 1(a) and Table 2(b) in paper in Table 1(b), respectively:
Table 0.D.1: mAP accuracy of BMDG using (a) our prototype mixing and (b) Mixup [mixup] setting for different numbers of part prototypes ($K$) and intermediate steps ($T$).
@@ 0.D.2 Step size and number of part prototypes:
0.D.3 Model efficiency:
Our BMDG proposed a method to extract alignable part-prototypes in feature extraction and then compute an attention embedding for the final representation features. In Table 0.D.2, we showed each component size and time complexity in the inference time compared to the baseline we used.
Table 0.D.2: Number parameters and floating-point operations at inference time for BMDG and all its sub-modules.
@@ 0.D.3 Model efficiency:
Appendix 0.E Visual Results
0.E.1 UMAP projections:
To show the effectiveness of BMDG, we randomly select 7 identities from the SYSU-MM01 dataset and project their feature representations using the UMAP method [mcinnes2018umap] for (a) Baseline, (b) one-step (prototypes without gradual training), and (c) BMDG. Visualization results (Figure 0.E.1) show that compared with the baseline and one-step approach, the feature representations learned with our BMDG method are well clustered according to their respective identity, showing a strong capacity to discriminate. BMDG is effective for learning robust and identity-aware features. Our BMDG method reduces this distance across modalities for each person and provides more separation among samples from different people.
@@ 0.E.1 UMAP projections:
Figure 0.E.1: Distributions of learned V and infrared features of 7 identities from SYSU-MM01 dataset for (a) the baseline, (b) one-step using part-prototypes, and (c) our BMDG method by UMAP [mcinnes2018umap]. Each color shows the identity.
Also, to show how the intermediate features gradually mix the modalities, we draw intermediate features for 6 steps in Figure 0.E.2. At the beginning of training, the features are based on modality while at step 6, the features are concentrated on each identity.
@@
Figure 0.E.2: Distributions of learned V and infrared features of 7 identities from SYSU-MM01 dataset in training for 6 steps at epochs of 10,30,50,70,90 and 160 respectively by UMAP [mcinnes2018umap]. Each color shows the identity. The intermediate features are drawn with lower opacity.
0.E.2 Domain shift:
To estimate the level of domain shift over data from V and I modalities, we measured the MMD distance for each training epoch. To this end, for each epoch, we selected 10 random images from 50 random identities and extracted the prototype and global features, then measured the MMD distance between the centers of those features for each modality as shown in Fig. 0.E.3(b). We report the normalized MMD distances between I and V features for our BMDG approach when compared with the baseline. Our method reduces this distance more than the y baseline. Thus, the results show that the intermediate domains improve the model robustness to a large multi-modal domain gap by gradually increasing the mix in prototypes over multiple steps.
@@ 0.E.2 Domain shift:
Figure 0.E.3: The mAP and domain shift (MMD distance) between I and V modalities over training epochs. (a) The learning curve of BMDG vs Baseline[all-survey]. (b) MMD distance over the center of multiple person’s infrared features to visible modality.
0.E.3 Part-prototype masking:
To show spatial information related to prototype features, we visualize the score map in the PM module (see Fig. 0.E.4). Our approach encodes prototype regions linked to similar body parts without considering person identity. Our model tries to find similar regions for each class of prototypes and then extracts ID-related information for that region. Therefore, BMDG is more robust for matching the same part features.

### v2

@@ 5 Conclusion
This paper introduces the BMDG framework for V-I ReID. It uses a novel prototype learning approach to learn the most discriminative and complementary set of prototypes from both modalities. These prototypes generate multiple intermediate domains between modalities by progressively mixing prototypes from each modality, reducing the domain gap. The effectiveness of BMDG is shown experimentally on several datasets, where it outperforms state-of-the-art methods for V-I person ReID. Results also highlight the impact of considering multiple intermediate steps and bidirectional training to improve accuracy. Moreover, BMDG can be integrated into other part-based V-I ReID methods, significantly improving their accuracy. A limitation of BMDG is its dependency on discriminant body parts.
Acknowledgements. This research was supported by the Natural Sciences and Eng. Research Council of Canada.
Appendix A Proofs
@@
A.1 Maximizing $\text{MI}(P^{1},\dots,P^{K};Y)$
@@
$\text{MI}(P^{1};Y)+\dots+\text{MI}(P^{K};Y),$ (A.1)
@@
$\text{MI}(X,Z;Y)=\text{MI}(X;Y)+\text{MI}(Z;Y)-\text{MI}(X;Z;Y)$ (A.5)
$MI(P^{1},\dots,P^{K};Y).\vspace{-.2cm}$ (A.6)
Figure A.1: Venn diagram of theoretic measures for three variables $X$, $Y$, and $Z$, represented by the lower left, upper, and lower right circles, respectively.
Theorem 1
Let $P^{1},\dots,P^{K}$ and $Y$ be random variables with domains $\mathcal{P}^{1},\dots,\mathcal{P}^{K}$ and $\mathcal{Y}$, respectively. Let every pair $P^{k}$ and $P^{q}$ ($k\neq q$) be independent. Then, maximizing $\text{MI}(P^{1},\dots,P^{K};Y)$ can be approximated by maximizing the sum of MI between each of $P^{k}$ to $Y$, $\sum_{k=1}^{K}{\text{MI}(P^{k};Y)}$.
Proof 1
@@
$\text{MI}(P^{1},\tilde{P}^{1};Y)=\underbrace{\text{MI}(P^{1};Y)}_{\alpha}+\underbrace{\text{MI}(\tilde{P}^{1};Y)}_{\beta}-\underbrace{\text{MI}(P^{1};\tilde{P}^{1};Y)}_{\gamma},$ (A.7)
@@
$\text{MI}(P^{1};\tilde{P}^{1};Y)\leq\text{MI}(P^{1};\tilde{P}^{1}),$ (A.8)
So by minimizing the right term of Eq. A.8, $\gamma$ is also minimized. We show that $\text{MI}(P^{1};\tilde{P}^{1})=0$ by expanding $\tilde{P}^{1}$ to $(P^{2},\tilde{P}^{2})$:
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=\text{MI}(P^{1};P^{2})+\text{MI}(P^{1};\tilde{P}^{2})-\text{MI}(P^{1};P^{2};\tilde{P}^{2}).$ (A.9)
@@
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=\text{MI}(P^{1};\tilde{P}^{2}).$ (A.10)
@@
$\text{MI}(P^{1};P^{2},\tilde{P}^{2})=0.$ (A.11)
To maximize $\beta$, we can rewrite and expand recursively in Eq. A.7 :
$\text{MI}(\tilde{P}^{1};Y)=\text{MI}(P^{2},\tilde{P}^{2};Y)=\underbrace{\text{MI}(P^{2};Y)}_{\text{maximizing}}+\underbrace{\text{MI}(\tilde{P}^{2};Y)}_{\text{expanding}}-\underbrace{\text{MI}(P^{2};\tilde{P}^{2};Y)}_{0}.$ (A.12)
@@
$\text{MI}(\tilde{P}^{k};Y)=\text{MI}(P^{k+1},\tilde{P}^{k+1};Y)=\text{MI}(P^{k};Y)+\dots+\text{MI}(P^{K};Y)\;\;\;\forall k\in\{0,\dots,K-1\}.$ (A.13)
@@
$\text{MI}(P^{1},\dots,P^{K};Y)=\sum_{k=1}^{K}{\text{MI}(P^{k};Y)},$ (A.14)
@@
A.2 Maximizing $\text{MI}(P^{k};Y)$
In Section 3, the MI between the representation of each prototype and the label of persons are maximized by minimizing cross-entropy loss (see Eq: 13 of the manuscript). This approximation is formulated as Proposition 1.
@@ A.2 Maximizing $\text{MI}(P^{k};Y)$
Let $P^{k}$ and $Y$ be random variables with domains $\mathcal{P}^{k}$ and $\mathcal{Y}$, respectively. Minimizing the conditional cross-entropy loss of predicted label $\hat{Y}$, denoted by $\mathcal{H}(Y;\hat{Y}|P^{k})$, is equivalent to maximizing the $\text{MI}(P^{k};Y)$
Proof 2
@@
$\text{MI}(P^{k};Y)=\underbrace{\mathcal{H}(Y)}_{\delta}-\underbrace{\mathcal{H}(Y|P^{k})}_{\xi}$ (A.15)
@@
$\mathcal{H}(Y;\hat{Y}|P^{k})=\mathcal{H}(Y|P^{k})+\underbrace{\mathcal{D}_{\text{KL}}(Y||\hat{Y}|P^{k})}_{\geq 0},$ (A.16)
@@
$\mathcal{H}(Y|P^{k})\leq\mathcal{H}(Y;\hat{Y}|P^{k}).$ (A.17)
Through the minimization of Eq. 13, training can naturally be decoupled in 2 steps. First, weights of the prototype module are fixed, and only the classifier parameters (i.e., weight $W^{k}$ of the fully connected layer) are minimized w.r.t. Eq. A.16. Through this step, $\mathcal{D}_{\text{KL}}(Y||\hat{Y}|P^{k})$ is minimized by adjusting $\hat{Y}$ while the $\mathcal{H}(Y|P^{k})$ does not change. In the second step, the prototype module’s weights are minimized w.r.t. $\mathcal{H}(Y|P^{k})$, while the classifier parameters $W^{k}$ are fixed.
Appendix B Additional Details on the proposed method
B.1 Training Algorithm
@@
B.2 Attentive Prototype Embedding
Details of the Attentive Prototype Embedding module are depicted in B.1.
Figure B.1: Attentive prototype embedding (APE) architecture.
Appendix C Additional Details on the Experimental Methodology
C.1 Datasets:
Research on cross-modal V-I ReID has extensively used the SYSU-MM01 [SYSU], RegDB [regDB], and recently published LLCM [LLCM] datasets. SYSU-MM01 is a large dataset containing more than 22K RGB and 11K IR images of 491 individuals captured with 4 RGB and 2 near-IR cameras, respectively. Of the 491 identities, 395 were dedicated to training, and 96 were dedicated to testing. Depending on the number of images in the gallery, the dataset has two evaluation modes: single-shot and multi-shot. RegDB contains 4,120 co-located V-I images of 412 individuals. Ten trial configurations randomly divide the dataset into two sets of 206 identities for training and testing. The tests are conducted in two ways – comparing I to V (query) and vice versa. An LLCM dataset consists of a large, low-light, cross-modality dataset that is divided into training and testing sets at a 2:1 ratio.
C.2 Experimental protocol:
We used a pre-trained ResNet50 [resnet] as the deep backbone model. Each batch contains 8 RGB and 8 IR images from 10 randomly selected identities. Each image input is resized to 288 by 144, then cropped and erased randomly, and filled with zero padding or mean pixels. ADAM optimizer with a linear warm-up strategy was used for the optimization process. We trained the model by 180 epochs, in which the initial learning rate is set to 0.0004 and is decreased by factors of 0.1 and 0.01 at 80 and 120 epochs, respectively. $K=6$, $T=4$, $\lambda_{\textit{f}}=0.1$, $\lambda_{\textit{v}}=0.05$, $\lambda_{\textit{p}}=0.2$ and $\lambda_{\textit{i}}=0.4$ are set based on the analyses shown in the ablation study in the main paper and in Section D.1. $\lambda_{\textit{eq}}=0.5$ is for all experiments.
C.3 Performance measures:
@@
Appendix D Additional Quantitative Results
D.1 Hyperparameter values:
This subsection analyzes the impact of $\lambda_{\textit{f}}$, $\lambda_{\textit{v}}$, $\lambda_{\textit{p}}$, and $\lambda_{\textit{i}}$ on V-I ReID accuracy. We initially set $\lambda_{\textit{v}}=0.01$, $\lambda_{\textit{p}}=0.05$, and $\lambda_{\textit{i}}=0.8$, experimenting with various values for $\lambda_{\textit{f}}$. As shown in Fig. D.1, accuracy improves with increasing $\lambda_{\textit{f}}$ until it reaches 0.1. Elevated $\lambda_{\textit{f}}$ enhances prototype diversity, boosting the discriminative ability of final features in the diverse space. However, excessively high values disperse prototype features in the feature space, diminishing discriminability and hindering accurate identification.
Similar trends are observed when $\lambda_{\textit{p}}=0.05$ and $\lambda_{\textit{i}}=0.8$, varying $\lambda_{\textit{v}}$ from 0.01 to 0.05, resulting in improved performance. Higher $\lambda_{\textit{v}}$ compresses prototype regions excessively, lacking sufficient ID-related information. Conversely, $\lambda_{\textit{i}}$ enhances the discriminative capabilities of prototypes in images. Balancing these factors, we find optimal values of 0.05 and 0.4 for $\lambda_{\textit{v}}$ and $\lambda_{\textit{i}}$, respectively. Additionally, based on experimentation, we set $\lambda_{\textit{p}}=0.2$ at the end of our analysis.
(a) $\lambda_{\textit{f}}$.
(b) $\lambda_{\textit{v}}$.
(c) $\lambda_{\textit{i}}$.
(d) $\lambda_{\textit{p}}$.
Figure D.1: Accuracy of the proposed BMDG over $\lambda_{\textit{f}}$, $\lambda_{\textit{v}}$, $\lambda_{\textit{i}}$, and $\lambda_{\textit{p}}$ values on SYSU-MM01 dataset in all-search and single-shot mode.
D.2 Step size and number of part prototypes:
In section 4.2, we discussed the step size and number of part prototypes based on Rank-1 accuracy. Here we report the mAP measurement for Table 4 in paper in Table 1(a) and Table LABEL:tab:part_step2 in paper in Table 1(b), respectively:
Table D.1: mAP accuracy of BMDG using (a) our prototype mixing and (b) Mixup [mixup] setting for different numbers of part prototypes ($K$) and intermediate steps ($T$).
@@ D.2 Step size and number of part prototypes:
D.3 Model efficiency:
Our BMDG proposed a method to extract alignable part-prototypes in feature extraction and then compute an attention embedding for the final representation features. In Table D.2, we showed each component size and time complexity in the inference time compared to the baseline we used.
Table D.2: Number parameters and floating-point operations at inference time for BMDG and all its sub-modules.
@@ D.3 Model efficiency:
Appendix E Visual Results
E.1 UMAP projections:
To show the effectiveness of BMDG, we randomly select 7 identities from the SYSU-MM01 dataset and project their feature representations using the UMAP method [mcinnes2018umap] for (a) Baseline, (b) one-step (prototypes without gradual training), and (c) BMDG. Visualization results (Figure E.1) show that compared with the baseline and one-step approach, the feature representations learned with our BMDG method are well clustered according to their respective identity, showing a strong capacity to discriminate. BMDG is effective for learning robust and identity-aware features. Our BMDG method reduces this distance across modalities for each person and provides more separation among samples from different people.
@@ E.1 UMAP projections:
Figure E.1: Distributions of learned V and infrared features of 7 identities from SYSU-MM01 dataset for (a) the baseline, (b) one-step using part-prototypes, and (c) our BMDG method by UMAP [mcinnes2018umap]. Each color shows the identity.
Also, to show how the intermediate features gradually mix the modalities, we draw intermediate features for 6 steps in Figure E.2. At the beginning of training, the features are based on modality while at step 6, the features are concentrated on each identity.
@@
Figure E.2: Distributions of learned V and infrared features of 7 identities from SYSU-MM01 dataset in training for 6 steps at epochs of 10,30,50,70,90 and 160 respectively by UMAP [mcinnes2018umap]. Each color shows the identity. The intermediate features are drawn with lower opacity.
E.2 Domain shift:
To estimate the level of domain shift over data from V and I modalities, we measured the MMD distance for each training epoch. To this end, for each epoch, we selected 10 random images from 50 random identities and extracted the prototype and global features, then measured the MMD distance between the centers of those features for each modality as shown in Fig. E.3(b). We report the normalized MMD distances between I and V features for our BMDG approach when compared with the baseline. Our method reduces this distance more than the y baseline. Thus, the results show that the intermediate domains improve the model robustness to a large multi-modal domain gap by gradually increasing the mix in prototypes over multiple steps.
@@ E.2 Domain shift:
Figure E.3: The mAP and domain shift (MMD distance) between I and V modalities over training epochs. (a) The learning curve of BMDG vs Baseline[all-survey]. (b) MMD distance over the center of multiple person’s infrared features to visible modality.
E.3 Part-prototype masking:
To show spatial information related to prototype features, we visualize the score map in the PM module (see Fig. E.4). Our approach encodes prototype regions linked to similar body parts without considering person identity. Our model tries to find similar regions for each class of prototypes and then extracts ID-related information for that region. Therefore, BMDG is more robust for matching the same part features.